OpenStudy (anonymous):

A gardener marks out a new section of lawn that is supposed to be a rectangle with sides of length 15 cm and 10cm respectively. to check he has marked out a rectangle he measures the length of the diagonal . how long should this be ?

OpenStudy (anonymous):

Ok have you ever used Pythagorean theorem before?

OpenStudy (anonymous):



OpenStudy (anonymous):

I forgot the scale

OpenStudy (anonymous):

Hm? You can solve it by plugging it into the formula c^2=a^2+b^2

OpenStudy (anonymous):

do you get how to do this?

OpenStudy (anonymous):

c*c=15*15+ 10*10?

OpenStudy (anonymous):

Yes that is right! so what is 15*15 and 10*10?

OpenStudy (anonymous):

325

OpenStudy (anonymous):

exactly so c^2=325, the opposite of square is the square root so you take the square root of both sides

OpenStudy (anonymous):

=18?

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

Thanks, Forgot the formula , has been a while to start the maths again

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

simplify 6^2+y^2-x^2+5+x-3y^3 ?

OpenStudy (anonymous):

is 5x^2+5+x+y^2 -3y^3 ?

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

\[\huge \hspace{-2pt}{\color{red}{\bigstar\sf Welcome~\to~Openstudy\bigstar}}\hspace{-250pt}{\color{violet}{\bigstar\sf Welcome~\to~Openstudy\bigstar}}\hspace{-250pt}{\color{blue}{\bigstar\sf Welcome~\to~Openstudy\bigstar}}\hspace{-250pt}{\color{green}{\bigstar\sf Welcome~\to~Openstudy\bigstar}}\hspace{-250pt}{\color{indigo}{\bigstar\sf Welcome~\to~Openstudy\bigstar}}\hspace{-250pt}{\color{yellow}{\bigstar\sf Welcome~\to~Openstudy\bigstar}}\hspace{-250pt}{\color{orange}{\bigstar\sf Welcome~\to~Openstudy\bigstar}}\hspace{-250pt}{\color{red}{\bigstar\sf Welcome~\to~Openstudy\bigstar}}\]